### v2

# G2P: Gaussian-to-Point Attribute Alignment for Boundary-Aware 3D Semantic Segmentation

Abstract
Semantic segmentation on point clouds is critical for 3D scene understanding. However, sparse and irregular point distributions provide limited appearance evidence, making geometry-only features insufficient to distinguish objects with similar shapes but distinct appearances (e.g., color, texture, material). We propose Gaussian-to-Point (G2P), which transfers appearance-aware attributes from 3D Gaussian Splatting to point clouds for more discriminative and appearance-consistent segmentation. Our G2P address the misalignment between optimized Gaussians and original point geometry by establishing point-wise correspondences. By leveraging Gaussian opacity attributes, we resolve the geometric ambiguity that limits existing models. Additionally, Gaussian scale attributes enable precise boundary localization in complex 3D scenes. Extensive experiments demonstrate that our approach achieves superior performance on standard benchmarks and shows significant improvements on geometrically challenging classes, all without any 2D or language supervision. Code will be released soon.
footnotetext: Equal contribution. Corresponding author.
1 Introduction
Figure 1: G2P augments Gaussian attributes to points while preserving geometry. (a) Point Cloud. Traditional point cloud representations lack sufficient appearance information to distinguish objects with similar geometry. As shown in the prediction, the existing model fails to segment coplanar windows cleanly and misclassifies the metallic refrigerator (yellow boxes), revealing geometric bias from insufficient appearance cues. (b) GS-Augmented Point Cloud (ours). Through our novel Gaussian-to-Point feature augmentation, G2P enriches points with Gaussian attributes $(\Sigma,\alpha)$, which encode rich appearance and geometric information. This enables successful segmentation of both challenging cases, resolving the geometric bias problem.
Point cloud segmentation is a fundamental task that enables comprehensive 3D scene understanding across diverse real-world applications. Despite strong progress [7, 48, 53, 26], the intrinsic sparsity and irregular sampling of point clouds force models to over-rely on coarse geometry, leading to geometric bias that confuses objects with similar shapes but distinct appearances [20]. For instance, background-adjacent objects (e.g., doors, windows, refrigerators) that are coplanar with walls or floors often become indistinguishable from their surroundings in sparse point clouds (Fig. 1(a)). This geometric bias highlights the need to align appearance cues for accurate discrimination.
Recent advances have improved geometric understanding of point clouds, but existing approaches still face two fundamental limitations. 1) Boundary ambiguity. Boundary-aware approaches [19, 47, 63] explicitly model edges, refining object boundaries. However, they remain limited to geometric reasoning and do not take advantage of the appearance information essential to address geometric bias [9, 15]. 2) Cross-modal misalignment. 2D-3D fusion approaches [46, 22, 6, 40, 16] inject rich image features but suffer from structural mismatches [15]. The discrete nature of point clouds makes wire-like structures hard to distinguish from flat surfaces [62], while projection introduces misalignment and occlusion-induced loss [28, 20]. These issues cause segmentation failures even for visually distinct regions, calling for a unified 3D representation that inherently encodes both geometry and appearance.
Among existing representations, 3D Gaussian Splatting (GS) [24] offers a promising direction that can address both geometric sparsity and appearance deficiency. Unlike point clouds with binary occupancy, Gaussian primitives possess continuous volumetric attributes that encode both geometric structure and appearance properties. Residing in the same 3D coordinate frame as points, these attributes can transfer with less spatial misalignment than 2D-3D fusion [60, 28, 62, 30]. However, optimization in GS often causes Gaussians to deviate from their initial positions [24, 33, 14, 43], creating a structural mismatch that prevents direct application of Gaussian attributes to point-level segmentation tasks.
To address this challenge, we introduce Gaussian-to-Point (G2P), which augments input points with Gaussian attributes while preserving their original geometry. As illustrated in Fig. 1(b), G2P creates appearance-aware point clouds by augmenting the original geometric information $(\mu^{p},c,n)$ with GS-derived attributes $(\Sigma,\alpha)$. This enables accurate segmentation of geometrically ambiguous objects that baseline methods tend to misclassify. Our G2P operates through three key components: (i) Gaussian-to-Point Feature Augmentation establishes precise correspondences via distance-based matching that accounts for anisotropic Gaussian ellipsoids; (ii) GS Primitives-guided Learning transfers view-consistent material cues to mitigate the point-level geometric bias; and (iii) Scale-based Boundary Extraction exploits anisotropic distributions to generate edge-aware pseudo-labels.
We conduct comprehensive experiments on multiple benchmark datasets. Our G2P consistently outperforms existing methods overall without any auxiliary supervision and achieves significant improvements on geometrically challenging classes. In summary, our main contributions are as follows:
  • We introduce G2P, the first approach that directly leverages aligned 3D Gaussian Splatting attributes for point cloud semantic segmentation.
  • We leverage appearance information in the native 3D space, bypassing cross-modal fusion limitations through Gaussian opacity-guided feature distillation.
  • We enhance segmentation accuracy via scale-based boundary extraction that leverages anisotropic Gaussian distributions.
2 Related Works
2.1 3D Point Cloud Segmentation
3D point cloud semantic segmentation has evolved from early per-point methods [29, 36, 51] to sophisticated architectures. PointNet++ [37] introduces hierarchical set abstraction for multi-scale learning, while MinkUNet [7] transforms point clouds into voxels and employs sparse convolutions to deliver robust and scalable segmentation. Recently, transformer-based architectures such as OctFormer [48] and Point Transformer v3 [53] have demonstrated state-of-the-art performance through efficient self-attention mechanisms.
Beyond supervised learning, self-supervised approaches like PointContrast [55] leverage contrastive learning, while Sonata [52] demonstrates effective pre-training on large-scale 3D datasets. Despite these advances, purely geometric approaches suffer from boundary ambiguity and geometric bias in sparse point clouds [56, 11], particularly where appearance cues become critical for distinguishing geometrically similar objects.
Boundary-Aware Approaches. Starting from Gong et al. [13], recent boundary-aware frameworks such as JSENet [19], CBL [47], and BFANet [63] explicitly learn boundary features to refine object boundaries in sparse point clouds [41]. These methods capture geometric discontinuities through dedicated boundary detection modules and edge-aware loss functions. However, they rely mainly on geometric cues and remain limited when segmentation requires appearance cues to address geometric bias arising from similar shapes but different materials.
Figure 2: Overview of the Gaussian-to-Point (G2P) learning approach. G2P comprises a preparation stage and a main training stage. In the preparation stage, (a) Gaussian-to-Point Feature Augmentation aligns 3D Gaussian ($\mu^{g},\Sigma,\alpha$) and point ($\mu^{p},c,n$), aggregating Gaussian attributes ($\alpha,S$), where $S$ represents the scale derived from the $\Sigma$, into augmented point features ($\alpha^{\prime},S^{\prime}$). (b) Scale-based Boundary Extraction uses the aggregated Gaussian scale magnitude to derive boundary pseudo-labels. An appearance encoder is pre-trained in a self-supervised manner using the augmented point representation ($\mu^{p},c,\alpha^{\prime}$). In the training stage, (c) GS Appearance Distillation transfers appearance cues from the pre-trained encoder to the 3D segmentation model, which takes point cloud inputs ($\mu^{p},c,n$). The model is trained under joint supervision from semantic labels, boundary pseudo-labels, and appearance distillation.
Cross-Modal 2D-3D Fusion. To address the appearance deficiency, cross-modal 2D-3D fusion approaches leverage rich visual features from RGB images. Multi-view methods such as VMVF [27], 3DMV [10], and MVPNet [22] aggregate features from multiple viewpoints through pooling or attention mechanisms. Subsequent studies [18, 40, 5, 20] further explore bidirectional projection and pre-trained backbone fusion. However, these approaches are hindered by point-to-pixel alignment errors in occluded regions, projection-induced information loss, and spatial misalignment between sparse 3D distributions and dense 2D grids [28, 15]. Consequently, 2D–3D fusion struggles to fully unify appearance and geometry within native 3D space [12, 57].
2.2 Gaussian Splatting for 3D Segmentation
With the advent of 3D Gaussian Splatting (GS) [24], recent works leverage GS for 3D segmentation and scene understanding [58, 45, 3], and even for open-vocabulary or open-world perception [54, 50, 23, 17, 31, 21, 61]. Beyond improvements in rendering quality, efficiency, and compression, Gaussians encode continuous volumetric geometry and appearance in 3D scenes. By using this potential, numerous methods [65, 38, 44, 67] lift CLIP [39]-/SAM [25]-derived features or masks into GS and optimize segmentation within the Gaussian representation. These methods focus primarily on view-consistent rendering with segmented Gaussians, which serves a different objective from point cloud segmentation. They also provide no explicit Gaussian-to-point correspondences to preserve the original geometric structure, and instead serve as an intermediate representation for rendering purposes.
UniPre3D [49] targets point-level tasks, but relies on pixel-wise rendering losses combined with cross-modal fusion to pretrain a backbone network, treating Gaussians as an auxiliary self-supervised signal rather than producing supervised point-level outputs. By contrast, our G2P transfers GS attributes directly to points via explicit Gaussian-to-point alignment under 3D supervision, without requiring recourse to 2D priors or rendering-based losses. Prior evidence shows that these Gaussian attributes encode both appearance and geometric properties [34, 30, 62]. Building on this insight, G2P exploits them as complementary appearance and boundary cues, coupling the appearance richness of Gaussians with the geometric stability of point clouds.
3 Method
3.1 Observation and Overview
In real-world 3D environments, numerous challenging classes lack clear geometric discriminability, which makes segmentation difficult. For example, doors and windows that are coplanar with walls, and appliances with reflective surfaces positioned adjacent to walls, are often geometrically indistinct from background structures. As shown in Fig. 1(a), the baseline model fails to distinguish between the refrigerator and the wall, misclassifying them as a single planar surface. Since these objects are primarily distinguishable by appearance cues such as color and texture, which are easily recognizable by human vision, learning representations solely based on geometry produces geometrically biased features.
To overcome the geometric bias of conventional methods, we propose Gaussian-to-Point (G2P), a unified learning approach that enriches point cloud representations with 3D Gaussian attributes. As illustrated in Fig. 2, G2P operates in two stages. In the Preparation Stage. (a) The Gaussian-to-Point Feature Augmentation module (Sec. 3.3) aligns Gaussians with points and augments geometric and appearance attributes, while (b) the Scale-based Boundary Extraction (Sec. 3.4) derives boundary pseudo-labels from the aggregated scale distribution. With the augmented point clouds, we train an appearance encoder. In the Main Training Stage. (c) GS Appearance Distillation (Sec. 3.5) employs the pre-trained appearance encoder to transfer appearance features to the point cloud segmentation network. This design integrates complementary geometric and appearance information within a unified 3D representation while preserving the original point geometry.
3.2 Preliminaries
3D GS is a view synthesis method that models a 3D scene as a set of anisotropic Gaussians and renders images through alpha blending. Each Gaussian is parameterized by centroid $\mu^{g}$, opacity $\alpha$, spherical harmonics $SH$, and a covariance matrix $\mathbf{\Sigma}$. The opacity $\alpha$ controls the blending contribution of each Gaussian along the ray, representing its visibility in the rendered view. The spatial shape of a Gaussian is defined by the covariance matrix $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ is a diagonal scale matrix, with the scale denoted as $S$ throughout. The scale $S$ determines the directional spread and geometric extent of the Gaussian in 3D space. We only use $S$ as a geometric cue and $\alpha$ as appearance information from Gaussian primitives. For additional analysis of Gaussian attributes, please refer to the Supplementary. These choices ensure that G2P focuses on informative attributes that are available for all scenes.
3D GS produces a set of optimized 3D Gaussians whose coordinates are altered during adaptive density optimization. As a result, the original geometric structure of the input point cloud is lost, and point-level semantic labels become unusable [24, 33, 14, 43]. The resulting Gaussian scene exhibits substantially more noise and structural artifacts, such as indistinct edges, when compared to the cleaner input point cloud, making direct use for segmentation unreliable.
3.3 Gaussian-to-Point Feature Augmentation
To address these fundamental limitations, we propose the Gaussian-to-Point feature augmentation approach. This approach maintains the original point cloud geometry and instead aligns each point with its nearest 3D Gaussians. Each point is then augmented with the associated Gaussian attributes, enabling semantic segmentation while preserving geometric fidelity. To augment the point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{N}$, where each point $p_{i}=(\mu_{i}^{p},c_{i},n_{i})\in\mathbb{R}^{9}$ consists of coordinates, color, and a normal vector, we transfer attributes from a set of 3D Gaussians $\mathcal{G}=\{g_{j}\}_{j=1}^{M}$. Both sets are first aligned in the same coordinate space. For each point $p_{i}$, we employ a two-stage process to identify its $k$ most relevant Gaussian neighbors. First, to efficiently narrow down candidates, we select all Gaussians whose centroids fall within a Euclidean radius $r$ of $p_{i}$. Second, within this candidate group, we compute the Mahalanobis distance [2] to each Gaussian. Unlike the Euclidean distance that assumes isotropic distributions, it incorporates each Gaussian’s anisotropic shape determined by its scale $S$ and rotation $R$, providing a more physically plausible measure of proximity. The final $k$ neighbors are selected based on the smallest Mahalanobis distances. Accordingly, the proposed weight parameter $w_{ij}$ for each neighbor $g_{j}$ is determined by the inverse of its Mahalanobis distance, normalized across the $k$ neighbors:
$w_{ij}=\frac{1/\sqrt{(\mu_{i}^{p}-\mu_{j}^{g})^{T}\Sigma_{j}^{-1}(\mu_{i}^{p}-\mu_{j}^{g})}}{\sum_{l=1}^{k}\left(1/\sqrt{(\mu_{i}^{p}-\mu_{l}^{g})^{T}\Sigma_{l}^{-1}(\mu_{i}^{p}-\mu_{l}^{g})}\right)}.$ (1)
These weights are first used to aggregate the scale $S_{j}^{g}\in\mathbb{R}^{3}$ and opacity $\alpha_{j}^{g}\in\mathbb{R}$ attributes from the $k$ neighbors. This yields the updated scale $S_{i}^{p}$ and weighted opacity $\alpha_{i}^{p}$ for each point $p_{i}$:
$S_{i}^{p}=\sum_{j=1}^{k}w_{ij}\cdot S_{j}^{g},\qquad\alpha_{i}^{p}=\sum_{j=1}^{k}w_{ij}\cdot\alpha_{j}^{g}.$ (2)
This augmentation extends each point $p_{i}$ to the form:
$p_{i}=(\mu^{p},c,n,S^{\prime},\alpha^{\prime})\in\mathbb{R}^{13}.$ (3)
As a result, the point cloud $\mathcal{P}$ is augmented with geometric $S^{\prime}$ and appearance $\alpha^{\prime}$ attributes from the Gaussians. This attribute-augmented point cloud is further employed to facilitate appearance-guided representation learning and to derive scale-based boundary pseudo-labels.
Figure 3: Gaussian Scale distribution. (a) Points after alignment, color-coded by aggregated Gaussian scale magnitude (magenta = small, yellow = large). Small scales concentrate at object boundaries. (b) Extracted boundary pseudo-labels by filtering out large-scale points.
3.4 Gaussian Scale-based Boundary Extraction
While geometric discontinuities can indicate boundaries, they fail for coplanar objects and thin structures that lack clear geometric separation. We instead exploit Gaussian scale attributes, which encode surface extent: small scales naturally concentrate at object edges and boundaries, while large scales cover smooth planar regions (Fig. 3).
Following the attribute augmentation, we leverage the newly acquired scale features $S_{i}^{\prime}$ from the augmented point cloud $\mathcal{P}^{\prime}$ to extract boundary pseudo-labels. The scale of 3D Gaussians reflects object geometry, with large-scale Gaussians typically covering broad planar regions. Based on this, we identify boundaries by filtering out points with large aggregated scales. Specifically, we first create an object-focused point cloud, $\mathcal{P}^{\prime}_{\text{obj}}$, by excluding points belonging to background classes (e.g., floor, wall) from $\mathcal{P}^{\prime}$. For each point $p_{i}^{\prime}$ with its augmented scale vector $S_{i}^{\prime}=(S_{x}^{\prime},S_{y}^{\prime},S_{z}^{\prime})$, we calculate its magnitude as the L2 norm:
$\|S_{i}^{\prime}\|_{2}=\sqrt{S_{x}^{\prime 2}+S_{y}^{\prime 2}+S_{z}^{\prime 2}}.$ (4)
Using $S_{i}^{\prime}$, we can now obtain the boundary pseudo-label subset $\mathcal{B}_{\text{scale}}$, which our scale-based approach identifies as the subset of $\mathcal{P}^{\prime}_{\text{obj}}$ whose scale magnitude is below a dynamically determined threshold $\tau_{\eta}$:
$\mathcal{B}_{\text{scale}}=\left\{p_{i}^{\prime}\in\mathcal{P}^{\prime}_{\text{obj}}\;\middle|\;\|S_{i}^{\prime}\|_{2}\leq\tau_{\eta}\right\},$ (5)
where the threshold $\tau_{\eta}$ is chosen to prune the top $\eta\%$ of points with the largest scale magnitudes, leaving the lower-scale points as boundary points.
To complement the geometric boundaries derived from scale, we also leverage semantic information. A point is identified as a semantic boundary candidate if any neighboring point within a local radius $r$ has a different semantic label, forming the set $\mathcal{B}_{\text{sem}}$. To yield a more robust set of boundary labels for subsequent training, we merge geometric and semantic cues through a logical OR (union) operation, forming the final pseudo-labels as $\mathcal{B}=\mathcal{B}_{\text{scale}}\cup\mathcal{B}_{\text{sem}}$.
3.5 Gaussian Attribute-guided Learning
Appearance Encoder Pre-training.
To preserve geometric consistency and prevent cross-modal misalignment, our approach takes point clouds as the sole input and receives appearance cues through feature distillation from a teacher encoder. To obtain such appearance representations, we repurpose the Sonata [52] architecture to pre-train this teacher encoder without pre-trained weights. While Sonata is a self-supervised framework that independently trains an encoder, in our setup, we configure this framework to learn appearance features. Instead of using the traditional representation $p_{i}=(\mu_{p},c,n)$ with 3D geometric normals $n\in\mathbb{R}^{3}$, we encode each point as $p_{i}=(\mu_{p},c,\alpha)$, where $\alpha\in\mathbb{R}$ denotes opacity, introducing view-consistent appearance cues in place of geometric information. The encoder is trained from scratch on the corresponding dataset (e.g., ScanNet v2 [8] or ScanNet++ [59]), ensuring dataset-specific pre-training for each benchmark. The pre-trained appearance encoder subsequently distills its representations into the 3D segmentation model that operates on point cloud $p_{i}=(\mu_{p},c,n)$.
Our approach incorporates boundary information into semantic segmentation by adopting the boundary-semantic (B-S) block from BFANet [63] and integrating it into the PT v3 [53] backbone. The B-S block disentangles the intermediate features into semantic and boundary representations using independent MLPs and enhances semantic features by attending to boundary-aware queries.
Table 1: Semantic segmentation on the ScanNet v2 validation set. Methods are grouped by training approach. denotes results reproduced by us. indicates external pre-training beyond ScanNet v2. Bold denotes the best in each column.
[…]
VMVF [27] ECCV’20 2D–3D Fusion 76.4
BPNet [18] CVPR’21 69.7
DeepViewAgg [40] CVPR’22 71.0
PonderV2 [66] TPAMI’25 77.0
UniPre3D [49] CVPR’25 77.6 85.0 92.1
ODIN (Swin-B) [20] CVPR’24 77.8
MinkUNet [7] CVPR’19 Geometric 72.3 80.2 90.0
OctFormer [48] TOG’23 74.3 82.6 90.9
PT v3 [53] CVPR’24 77.0 84.3 92.1
BFANet [63] CVPR’25 77.3 84.1 92.2
G2P (Ours) GS-guided 78.4 85.2 92.5
Training Losses.
Finally, the 3D segmentation model is trained using the proposed three supervisory signals: appearance, semantic, and boundary. For the appearance feature signal, we employ an additional mapping MLP $\phi$, which takes the output features of the main point encoder, $f_{i}^{p}$, as input. We then distill knowledge from the features of the appearance encoder $f_{i}^{a}$ by minimizing the cosine similarity loss, $\mathcal{L}_{\text{distill}}$ (the appearance distillation loss):
$\mathcal{L}_{\text{distill}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{\phi(f_{i}^{p})\cdot f_{i}^{a}}{\|\phi(f_{i}^{p})\|_{2}\|f_{i}^{a}\|_{2}}\right).$ (6)
Semantic supervision loss, $\mathcal{L}_{\text{sem}}$, is applied via a combined loss of cross-entropy and Lovász-softmax [1], while boundary supervision loss, $\mathcal{L}_{\text{bou}}$, is provided using a combination of binary cross-entropy and Dice loss [35].
For these losses, we denote the ground truth semantic labels as $y^{g}$, boundary pseudo-labels as $b^{g}$, and the corresponding model predictions as $\hat{p}$ and $\hat{b}$, respectively. Here, the loss functions are defined as:
[…]
Table 2: Class-wise IoU comparison on all ScanNet v2 categories. Red and blue denote the best and second-best IoU, respectively. $\dagger$ reproduced by us. Avg.: Average IoU for each subset. Refrig.: Refrigerator. ShwrCurt.: Shower Curtain.
Method Geometrically Distinguishable Classes Geometrically Challenging Classes
Wall Floor Cabinet Bed Chair Sofa Table Bookshelf Avg. Door Window Picture Curtain Refrig. ShwrCurt. Avg.
MinkUNet [7] 87.1 96.7 67.0 83.4 92.6 84.5 77.0 81.4 83.7 70.4 65.0 37.8 77.2 64.0 67.2 63.6
OctFormer [48] 86.1 95.7 70.1 82.7 91.9 83.6 74.1 81.3 83.2 67.7 67.2 34.0 77.4 65.5 65.8 62.9
PT v3 [53] 87.7 95.8 73.0 83.2 93.1 83.7 78.9 80.6 84.5 74.6 72.8 41.8 77.9 64.9 68.9 66.8
UniPre3D [49] 87.5 95.5 72.6 84.4 92.7 84.9 80.0 82.3 85.0 74.6 72.6 38.5 79.0 70.3 76.3 68.6
BFANet [63] 87.2 95.7 72.3 84.7 92.5 83.0 81.2 80.6 84.7 74.0 73.5 38.5 77.5 69.3 70.8 67.3
G2P (Ours) 87.6 95.8 73.3 87.8 93.0 84.6 79.7 84.6 85.8 76.8 73.9 39.4 78.0 70.9 76.1 69.2
[…]
The final loss, $\mathcal{L}_{\text{total}}$, is a weighted sum of the three terms, where $\lambda_{b}$ and $\lambda_{d}$ are weights balancing the boundary and distillation losses, respectively:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sem}}+\lambda_{b}\mathcal{L}_{\text{bou}}+\lambda_{d}\mathcal{L}_{\text{distill}}.$ (9)
4 Experiments
4.1 Experiment Setting
Dataset. We evaluate our approach on the ScanNet v2 [8, 42], ScanNet++ [59], and Matterport3D [4] datasets. ScanNet v2, ScanNet200, ScanNet++, and Matterport3D contain 20, 200, 100, and 21 semantic classes, respectively. We report results on the validation splits of all datasets. For 3D Gaussian representations, we adopt the SceneSplat-7K [30] dataset, where each scene is reconstructed with approximately 1.5M Gaussian primitives. SceneSplat-7K provides GS reconstructions for indoor scenes, but outdoor environments are not available in the dataset.
Table 3: Semantic segmentation on the ScanNet200 validation set. $\dagger$ denotes reproduced results. indicates external pre-training beyond ScanNet v2. Bold denotes the best in each column.
[…]
MinkUNet [7] CVPR’19 25.0 32.9 80.4
PointContrast [55] ECCV’20 26.2
OneFormer3D [26] CVPR’24 30.1
PonderV2 [66] TPAMI’25 32.3 - -
OctFormer [48] TOG’23 32.6 40.1 83.0
PT v3 [53] CVPR’24 34.5 45.4 82.7
UniPre3D [49] CVPR’25 36.0 46.2 83.7
G2P (Ours) 36.6 46.0 83.8
Evaluation Metric. Following prior works [53, 63], we adopt three standard metrics for evaluation: mean Intersection over Union (mIoU), mean Accuracy (mAcc), and overall Accuracy (OA), measuring the percentage of correctly classified points across the entire dataset.
Implementation details. Our main training is conducted on a single NVIDIA RTX 3090 GPU for 800 epochs. The appearance encoder pre-training, following Sonata [52], is performed separately on an NVIDIA A6000 GPU for 400 epochs per dataset from scratch. We adopt Point Transformer v3 (PT v3) [53] as the backbone and incorporate the B-S block from BFANet [63]. The training batch size is set to 4, and we adopt Adam [32] as the optimizer with an initial learning rate of 0.003. All other settings follow PT v3, and detailed architectural configurations are reported in the Supplementary. For boundary pseudo-label extraction, we set the scale-based trimming ratio $\tau_{\eta}=0.7$. The radius $r$ for semantic boundary calculation is set to 4 cm. Loss weights are balanced as $\lambda_{d}$ = 0.4, $\lambda_{b}$ = 0.9.
Figure 4: Qualitative comparison on geometrically challenging classes in ScanNet v2. Yellow boxes highlight regions with differing results between models. Boxes with stars ($\star$) indicate category confusion, where baselines misclassify objects (shown in blue) due to geometric bias (e.g., door $\rightarrow$ wall, refrigerator $\rightarrow$ cabinet). Boxes without stars indicate incomplete segmentation, where baselines fail to cover the full extent of thin or coplanar structures. G2P mitigates both failure modes through opacity-guided appearance distillation and scale-based boundary extraction. Corresponding regions in Ground Truth (bottom row) are marked for reference. Best viewed in color with zoom.
4.2 Quantitative Comparison
Results on ScanNet v2.
We compare G2P with geometric methods and 2D-3D fusion approaches on the ScanNet v2 validation set. Tab. 1 reports the results. G2P exceeds PT v3 by +1.4 mIoU and BFANet [63] by +1.1 under identical settings. G2P outperforms all geometric and 2D–3D fusion baselines without external pre-training. For reference, UniPre3D [49] and ODIN [20] both use external pre-training beyond ScanNet v2.
Class-wise analysis. To better understand where G2P excels, we categorize the 20 classes in ScanNet v2 into two groups based on geometric discriminability. Geometrically challenging classes include objects that are difficult to distinguish by geometry alone and are often coplanar with walls or exhibit reflective surfaces. Geometrically distinguishable classes include objects with clear structural features. As shown in Tab. 2, our G2P demonstrates superior performance on geometrically challenging classes with an average IoU of 69.2%. The largest class-level improvements occur on refrigerator (70.9 IoU, +6.0 vs. PT v3’s 64.9) and shower curtain (+7.2).
Table 4: Evaluation on ScanNet200 Hidden Test Set. We report mIoU, Head, Common, and Tail IoU on the official ScanNet200 benchmark (as of November 2025). Red and blue denote the best and second-best in each metric.
Method Venue mIoU Head Common Tail
MinkUNet [7] CVPR’19 25.3 46.3 15.4 10.6
OctFormer [48] TOG’23 32.6 53.9 26.5 13.1
CeCo [64] CVPR’23 34.0 55.1 24.7 18.1
PonderV2 [66] CVPR’24 34.6 55.2 27.0 17.5
BFANet [63] CVPR’25 36.0 55.3 29.3 19.3
G2P (Ours) 35.5 55.3 28.6 18.7
Results on ScanNet200. We further evaluate on the ScanNet200 [42] validation set, which expands the label space from 20 to 200 fine-grained categories (Tab. 3). G2P attains 36.6 mIoU, 46.0 mAcc, and 83.8 OA, improving over our reproduced PT v3 [53] baseline by +2.1 mIoU and remaining competitive with UniPre3D [49]. We do not report BFANet [63] on ScanNet200 due to the lack of official implementation for this dataset.
Table 5: Module ablation on ScanNet v2. (a) Baseline uses pure geometric features. (b) adds boundary guidance to improve localization on coplanar and thin structures. (c) adds appearance distillation to stabilize semantic consistency in category interiors.
Method Components Performance
[…]
Guidance Distillation
(a) 77.0 92.1
(b) 77.8 92.1
(c) 77.8 92.0
G2P (Ours) 78.4 92.5
We further evaluate our method on the ScanNet200 hidden test set (Tab. 4). Following BFANet, we compare with methods that do not use large-scale pre-training beyond ScanNet v2. G2P achieves competitive performance with 35.5 mIoU, closely approaching BFANet’s 36.0 while matching it on Head classes and maintaining strong performance across Common and Tail categories.
4.3 Qualitative Comparisons
Qualitative results on challenging classes. Fig. 4 compares G2P with PT v3 [53] and BFANet [63] on challenging ScanNet v2 classes. Baselines exhibit two primary failure modes, highlighted by yellow boxes. Boxes with stars ($\star$) mark category confusion: in columns (a) door and (e) refrigerator, the baselines misclassify the objects as wall or cabinet due to coplanar geometry and similar textures. G2P correctly segments these objects. Boxes without stars mark incomplete segmentation: in columns (b) window, (c) picture, (d) curtain, and (f) shower curtain, baselines fail to cover the full extent of thin structures. G2P produces sharper, more complete masks for these objects. These trends align with Tab. 2, where G2P achieves substantial gains on refrigerator (improvements of 6.0% in IoU vs. PT v3), confirming that G2P effectively mitigates both failure modes.
4.4 Ablation Studies
Module ablation. We conduct ablation studies to validate the contribution of each component (Tab. 5). Both boundary guidance and appearance distillation independently improve mIoU, with their combination achieving the highest gain. This confirms that boundary and appearance cues from 3D Gaussians provide complementary benefits. Each component alone produces competitive results, and combining them yields the most reliable segmentation by reducing boundary leakage and false merges.
Results on Additional Benchmarks. To further validate G2P’s effectiveness, we evaluate on the ScanNet++ [59] and Matterport3D [4] validation sets. As shown in Tab. 6, G2P generalizes well across datasets, achieving 48.7 and 55.9 mIoU on ScanNet++ and Matterport3D, outperforming PT v3 by +0.8 and +0.4.
Table 6: Results on ScanNet++ and Matterport3D. All models are trained and evaluated within each dataset. Bold indicates the best in each column. $\dagger$ denotes reproduced results.
Method ScanNet++ (val) Matterport3D (val)
[…]
PT v3 [53] 47.9 59.1 90.0 55.5 64.9 82.1
G2P (Ours) 48.7 60.1 90.0 55.9 66.4 82.2
Appearance transfer. Tab. 7 evaluates strategies for leveraging Gaussian appearance features. Using raw Gaussian coordinates proves unreliable for point-level segmentation (73.1 mIoU). In contrast, Points with Opacity improves performance to 77.3 mIoU, demonstrating that Gaussian appearance carries useful cues for segmentation. Our distilled variant reaches the highest mIoU of 77.8 and removes the need for Gaussian features at inference, in contrast to Points with Opacity method.
Table 7: Ablation on strategies for Gaussian primitive (opacity) Learning. We compare: (i) Raw 3D Gaussians, which use Gaussian primitives as direct training inputs; (ii) Points with Opacity, which also provide point features and their Gaussian opacity directly to the model; and (iii) G2P (Ours), which distills Gaussian appearance to the segmentation model. Here, $c^{\prime}$ and $n^{\prime}$ denote Gaussian color and normal features aggregated from neighboring points based on Euclidean distance.
[…]
Raw 3D Gaussians Direct input $x^{g},c^{\prime},n^{\prime},\alpha$ 73.1
Points with Opacity Direct input $x^{p},c,n,\alpha^{\prime}$ 77.3
G2P (Ours) Distillation $x^{p},c,n,\alpha^{\prime}$ 77.8
Correspondence metric. Gaussian-to-point correspondence depends on the distance metric and the neighborhood size $k$. Euclidean distance ignores the anisotropy of splats and yields unstable matches near object boundaries. Mahalanobis distance leverages the covariance $\boldsymbol{\Sigma}$ to form tighter, more reliable matches. As shown in Tab. 8, a moderate neighborhood ($k=20$) offers the best trade-off (78.4 mIoU, 92.5 OA). Smaller $k$ is noisy, whereas larger $k$ oversmooths labels. These results support covariance-aware matching as the default in G2P.
Inference efficiency. Tab. 9 evaluates computational efficiency. Training G2P incurs additional overhead due to appearance distillation: training latency increases from 132ms to 220ms (+67%) and memory usage from 5.6GB to 7.3GB (+30%). However, inference overhead remains moderate: parameters increase by only 0.2M (+0.4%), latency increases from 79ms to 96ms (+21%), and memory usage increases from 1.9GB to 3.3GB (+74%).
Table 8: Ablation on distance metric and neighborhood size for Gaussian-to-point feature augmentation (ScanNet v2). This experiment examines how different correspondence settings, including distance metric and the neighborhood size $k$, influence the alignment of Gaussian primitives to points during feature augmentation, which subsequently affects appearance encoder pre-training and boundary extraction.
[…]
Euclidean Distance 77.0 84.4 91.8
Mahalanobis (k=10k\!=\!10) 77.4 85.3 92.0
Mahalanobis (k=20k\!=\!20) 78.4 85.2 92.5
Mahalanobis (k=30k\!=\!30) 77.9 85.3 92.1
Table 9: Model efficiency on ScanNet v2. Comparison of training and inference efficiency between PT v3 and G2P. $\dagger$ denotes reproduced results.
[…]
Latency Memory Latency Memory
MinkUNet [7] 39.2M 71ms 1.6G 29ms 1.4G 72.3
OctFormer [48] 44.0M 259ms 4.2G 94ms 4.4G 74.3
PT v3 [53] 46.2M 132ms 5.6G 79ms 1.9G 77.0
G2P (Ours) 46.4M 220ms 7.3G 96ms 3.3G 78.4
5 Conclusion
We presented G2P, a novel approach for point cloud semantic segmentation that integrates 3DGS to unify geometric and appearance cues in 3D space. Our three-component approach, comprising Gaussian-to-point feature augmentation, opacity-guided appearance learning, and scale-based boundary extraction, addresses the fundamental geometric bias. By augmenting point features with Gaussian attributes, G2P enhances semantic segmentation on point clouds, particularly for objects that are geometrically ambiguous but have distinctive appearances. Extensive experiments on standard benchmarks demonstrate state-of-the-art performance and strong generalization across datasets. Furthermore, we conduct ablations to validate the effectiveness of GS-guided feature augmentation, establishing its practical value for 3D scene understanding.